# Fixing the Double Penalty in Data-Driven Weather Forecasting Through a Modified Spherical Harmonic Loss Function

### v1

@@ B.3 Details of the lagged ensemble verification
$\operatorname{eRMSE}(x,y)$ $=\left(\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\sum_{i}\sum_{j}\mathrm{d}A(i,j)\left(\frac{1}{N_{e}}\sum_{k=1}^{N_{e}}x_{k}(i,j)-y(i,j)\right)^{2}\right)^{1/2},\text{ and}$ (14)
$\operatorname{SER}(x,y)$ $=\left(\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\sum_{i}\sum_{j}\mathrm{d}A(i,j)\frac{\sum_{k=1}^{N_{e}}(x_{k}(i,j)-\bar{x}(i,j))^{2}}{(\sum_{k=1}^{N_{e}}x_{k}(i,j)-y(i,j))^{2}}\right)^{1/2},$ (15)
where $\bar{x}(i,j)$ is the ensemble mean at the $(i,j)$ gridpoint.
Figure 10: CRPS skill score (improvement), measured as the relative difference between the CRPS of the 12-step fine-tuned model and the CRPS of the control model, for a selection of variables and lead times. Orange up-arrows show where the fine-tuned model performs better, blue down-arrows show where the control model performs better. Hollow arrows represent a difference of less than 1%, and differences of 2% or larger are marked. Hollow circles mark values that are not statistically significant at the 90% level.
Figure 11: As figure 10, for ensemble root mean squared error.
Figures 10 and 11 show the CRPS and eRMSE skill scores respectively of the lagged ensemble generated with AMSE AR12 compared to the lagged ensemble of the control model for the geopotential (z), temperature (t), specific humidity (q), and u-component of wind (u) at several elevations and for the mean sea level pressure (msl), 2-meter temperature (2t), u-component of 10m wind (10u), and 6h-accumulated precipitation (tp) at the surface.
@@
The eRMSE skill chart should be interpreted with caution. The scores of (13)–(15) were developed for the case of an ideal ensemble, where members are statistically indistinguishable from each other and equally accurate in expectation. This is not really the case for a lagged ensemble, where the shorter-duration members should be noticeably more accurate than longer-duration members and an ideal aggregation would separately weight each term. This is not done by Brenowitz et al. [2024] for simplicity and to avoid free parameters, but we believe that the early lead-time smoothing in the control model makes equal-weighting more optimal for its lagged ensemble than for the lagged ensemble of AMSE AR12.

### v2

@@ B.4 Details of the Lagged Ensemble Verification
$\operatorname{eRMSE}(x,y)$ $=\left(\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\sum_{i,j}\mathrm{d}A(i,j)(\bar{x}(i,j)-y(i,j))^{2}\right)^{1/2},\text{ and}$ (15)
$\operatorname{SER}(x,y)$ $=\left(\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\frac{1}{N_{e}-1}\frac{\sum_{i,j}\mathrm{d}A(i,j)\sum_{k=1}^{N_{e}}(x_{k}(i,j)-\bar{x}(i,j))^{2}}{\sum_{i,j}\mathrm{d}A(i,j)(\bar{x}(i,j)-y(i,j))^{2}}\right)^{1/2},$ (16)
where $\bar{x}(i,j)=N_{e}^{-1}\sum_{k}x_{k}(i,j)$ is the ensemble mean at the $(i,j)$ gridpoint.
Figure 11: CRPS skill score (improvement), measured as the relative difference between the CRPS (14) of the 12-step fine-tuned model and the CRPS of the control model, for a selection of variables and lead times. Orange up-arrows show where the fine-tuned model performs better, blue down-arrows show where the control model performs better. Hollow arrows represent a difference of less than 1%, and differences of 2% or larger are marked. Hollow circles mark values that are not statistically significant at the 90% level.
Figure 12: As figure 11, for ensemble root mean squared error (15).
Figures 11 and 12 show the CRPS and eRMSE skill scores respectively of the lagged ensemble generated with AMSE AR12 compared to the lagged ensemble of the control model for the geopotential (z), temperature (t), specific humidity (q), and u-component of wind (u) at several elevations and for the mean sea level pressure (msl), 2-meter temperature (2t), u-component of 10m wind (10u), and 6h-accumulated precipitation (tp) at the surface.
@@
The eRMSE skill chart should be interpreted with caution. The scores of (14)–(16) were developed for the case of an ideal ensemble, where members are statistically indistinguishable from each other and equally accurate in expectation. This is not really the case for a lagged ensemble, where the shorter-duration members should be noticeably more accurate than longer-duration members and an ideal aggregation would separately weight each term. This is not done by Brenowitz et al. (2025) for simplicity and to avoid free parameters, but we believe that the early lead-time smoothing in the control model makes equal-weighting more optimal for its lagged ensemble than for the lagged ensemble of AMSE AR12.
@@
B.4.1 Unbiased Ensemble Root Mean Squared Error
Figure 13: As figures 11 and 12, for the unbiased ensemble root mean squared error (18).
The eRMSE formula of (15) is a biased estimator of the true ensemble mean error, overestimating the error in proportion to the ensemble (sample) spread when the ensemble size is finite.
Consider $N_{e}$ different realizations $({x_{i}})$ of a single variable drawn from $\operatorname{\mathcal{N}}(\mu,\sigma^{2})$ when the ground-truth value is 00. Applying (15) to this gives:
$\operatorname{\mathbb{E}}{(\mathrm{eMSE}(x,0))}=\operatorname{\mathbb{E}}\left(\left(\frac{1}{N_{e}}\sum_{i}x_{i}\right)^{2}\right)=\mu^{2}+\frac{\sigma^{2}}{N_{e}},$ (17)
which overestimates the true ensemble mean squared error. This overestimate is more severe for small ensembles such as the lagged ensemble configuration of section 3.2, where the ensemble size cannot be easily increased.
Leutbecher & Palmer (2008) proposes correcting this overestimate by subtracting the standard error term to give an unbiased estimator of the ensemble mean squared error with a finite sample size. In the notation of (15), the corresponding root mean squared formula becomes:
$\operatorname{ub\_eRMSE}(x,y)=\Bigg{(}\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}$ $\sum_{i,j}\mathrm{d}A(i,j)\bigg{(}\Big{(}\frac{1}{N_{e}}\sum_{k=1}^{N_{e}}x_{k}(i,j)-y(i,j)\Big{)}^{2}-$
$\frac{1}{N_{e}(N_{e}-1)}\sum_{k=1}^{N_{e}}(x_{k}(i,j)-\bar{x}(i,j))^{2}\bigg{)}\Bigg{)}^{1/2},$ (18)
which performs this correction pointwise on the grid before computing the spatial average and taking the square root.
Implementing this adjustment slightly improves the scores of the AMSE-tuned model compared to the control model, and the corresponding “scorecard” is shown in figure 13.
B.5 Ablation Studies
Figure 14: As figure 3, for the comparison models of section B.5. Only the model trained with the AMSE error function retains sharpness to fine scales.
Table 2: Cumulative ranked probability scores for the models fine-tuned in this study in the lagged ensemble configuration described in section 3.2, for the “headline” variables and levels in figure 1 of Rasp et al. (2024). Lower is better; the best score is bolded and the second-place score is italicized.
Model z 500hPa ($\mathrm{m}^{2}\mathrm{s}^{-2}$) t 850hPa ($\mathrm{K}$) q 700hPa ($\mathrm{g}\cdot\mathrm{kg}^{-1}$) u 850hPa ($\mathrm{m}\cdot\mathrm{s}^{-1}$)
2.5d 5.0d 7.5d 2.5d 5.0d 7.5d 2.5d 5.0d 7.5d 2.5d 5.0d 7.5d
Control 31.038 84.315 162.481 0.428 0.691 1.003 0.357 0.523 0.652 0.823 1.340 1.904
MSE AR12 31.285 82.100 155.703 0.419 0.664 0.949 0.356 0.526 0.652 0.819 1.335 1.886
MAE AR12 29.969 80.621 155.361 0.410 0.654 0.947 0.340 0.499 0.624 0.811 1.313 1.859
AMSE AR1 33.720 94.703 186.202 0.422 0.721 1.078 0.354 0.558 0.721 0.863 1.485 2.115
AMSE AR12 30.565 80.469 153.267 0.418 0.653 0.935 0.347 0.510 0.634 0.832 1.341 1.882
To ensure that the results of this study are not simply an artifact of increasing the model’s overall training time, we compare against two additional fine-tunings:
  1. 1.
    MSE AR12 implements the fine-tuning schedule of table 1 with the unmodified mean squared error loss function, as with GraphCast’s principal training.
  2. 2.
    MAE AR12 implements the fine-tuning schedule with a mean absolute error loss function, preserving the per-variable and per-level weightings of error.
Figure 14 shows the aggregated per-wavenumber performance of these models, and table 2 evaluates their CRPS for a selection of variables, levels, and lead times in the lagged ensemble configuration.
Both models still show excessive smoothing of fine scales, but training with mean absolute error moderately improves sharpness at the medium scales (wavenumbers 32–200 for longer lead times, corresponding to length scales of 1250–250 kilometers).
The excessive smoothing of the MSE-trained model is expected from section 2.1, but that argument does not directly apply to the mean absolute error loss function. However, we can still understand this behaviour intuitively. A model that is optimal under the mean absolute error predicts the mean of a distribution, and at longer lead times fine scales are less predictable than coarser scales. Therefore, the prediction of the median future should be smoother than its realization.
Even the moderate improvement to sharpness for the MAE-trained model results in improvements to the CRPS of the lagged ensemble, as shown in table 2.